OpenStudy (anonymous):

Sandra throws an object into the air with an initial vertical velocity of 38 feet per second, from a platform that is 30 feet above the ground. How long will it take the object to hit the ground? A. 2 seconds B. 4 seconds C. 1 second D. 3 seconds D, right? @Dscdago @IMStuck

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

 using : \(\sf y_f=y_i+v_y\Delta t +\dfrac{1}{2}a_y(\Delta t)^2\) \(\sf 0=30~ft+(38~f/s)\Delta t +\dfrac{1}{2}(-32~ft/s^2)(\Delta t)^2\) then \(\sf \Delta t=3~s\) so yeah, youre right

OpenStudy (imstuck):

Yes, it's 3.